United States Rule math


A partial payment is made on the date indicated. Use the united states rule to determine the balance due on the note at the date of maturity? Assume the year is not a leap year. Principle- $4000, Rate 4%, Effective date May ,Partial payment amount- $2000,Date June 1,Maturity date July 1
Answer #1
Interest on principal from April 1 to May 1 i.e. for 31 days
Interest = 4000*4%*31/365 13.59
Principal due after partpayment = 4000 + 13.59 - 2000 2013.59
Interest from May 1 to June 1 i.e. for 31 days
Interest = 2013.59*4%*31/365 6.84
Amount due on maturity = 2013.59 + 6.84 2020.43